# Modeling fibrous tissue in vascular fluid-structure interaction: a morphology-based pipeline and biomechanical significance

### v1

@@ 2.3 Constitutive model for arterial wall
In this section, we introduce the arterial wall constitutive models. The arterial wall is modeled as a slightly compressible or fully incompressible hyperelastic material. Given the referential configuration $\Omega_{0}\subset\mathbb{R}^{3}$, its motion at time $t$ is given by an orientation-preserving one-to-one mapping $\bm{\varphi}_{t}(\cdot)$ that maps the body to the current configuration $\Omega_{t}=\bm{\varphi}_{t}(\Omega_{0})$. The kinematics is locally described by the deformation gradient $\bm{F}:=\partial\bm{\varphi}_{t}/\partial\bm{X}$ for $\bm{X}\in\Omega_{0}$. Its determinant is known as Jacobian and denoted by $J$. The right Cauchy-Green deformation tensor is given by $\bm{C}:=\bm{F}^{T}\bm{F}$. On the referential configuration, a scalar field $P$ is introduced to represent the thermodynamic pressure. The material behavior is characterized by the free energy assumed of the form,
@@
where $G$ represents the Gibbs free energy. It is a scalar-valued function of $\bm{C}$ and the pressure $P$ on the referential configuration. It is additively decomposed into the contributions of $G_{\mathrm{d}}$ and $G_{\mathrm{v}}$. The deformation is often decoupled into isochoric and volumetric parts following the decomposition of the Flory type, i.e., $\bm{F}=J^{1/3}\tilde{\bm{F}}$, and correspondingly the unimodular right Cauchy-Green deformation tensor is defined as $\tilde{\bm{C}}:=\tilde{\bm{F}}^{T}\tilde{\bm{F}}$. If the energy function $G_{\mathrm{a}}$ depends on the deformation tensor through $\tilde{\bm{C}}$, it can be interpreted as the energy associated with isochoric deformation, and $G_{\mathrm{b}}$ is the bulk or volumetric free energy. Through a Coleman-Noll-type analysis, we may obtain the constitutive relations for the second Piola-Kirchhoff stress $\bm{S}$, the Cauchy stress $\bm{\sigma}$, and the density $\rho$ based on the free energy $G$ as
@@
In the theory constructed based on the Gibbs free energy, the pressure $P$ acts as an primitive variable, rendering a mixed formulation suitable for (nearly-)incompressible materials [48]. If the energy is taken simply as $G_{\mathrm{b}}=P$, the one has $\rho=\rho_{0}$ and the material is fully incompressible. In this work, we adopt the following form for the volumetric energy,
@@
where $K$ is the bulk modulus. This volumetric energy is related to the volumetric energy $K/4(J^{2}-1-2\ln(J))$ through a Legendre transformation [48]. The free energy $G_{\mathrm{d}}$ is directly inherited from the Helmholtz free energy, and a specific functional form of it completes the definition of the constitutive model.
@@
where $\tilde{I}_{1}:=\mathrm{tr}\tilde{\bm{C}}$ is the first principal invariant of $\tilde{\bm{C}}$, and the parameter $\mu$ is the shear modulus. This model enjoys a compact energy form with a single material parameter and can be rigorously derived based on the microscopic structure using the Gaussian distribution. Despite its wide usage in vascular FSI analysis, we need to point out that it was designed for vulcanized rubber and its predictive capability is limited to moderate deformation states [81].
@@
and the latter two describe the stretch along the fiber directions. The parameters $\mu$ and $k_{1}$ are stress‐like coefficients representing the stiffness of the ground matrix and collagen fibers; $k_{2}$ is a dimensionless parameter; the unit vectors $\bm{a}_{0}$ and $\bm{b}_{0}$ represent the mean orientation of the embedded fibers; the parameter $\kappa\in[0,1/3]$ characterizes the fiber dispersion. In the model (2.5), the first term represents the potential energy due to the ground matrix, which is treated as an isotropic neo-Hookean material. The summed exponential forms in (2.5) represent the anisotropic free energy contributions due to the two families of fibers.
In this work, we consider the two family of fibers disposed in a symmetric manner with respect to the axial direction. Therefore, the fiber directions $\mathbf{a}_{0}(\bm{X})$ and $\mathbf{b}_{0}(\bm{X})$ at a material point $\bm{X}\in\Omega_{0}$ can be defined by the local basis vector $\{\bm{e}_{\mathrm{r}}(\bm{X}),\bm{e}_{\mathrm{c}}(\bm{X}),\bm{e}_{\mathrm{a}}(\bm{X})\}$ as follow,
@@
where $\theta\in\left[0,\pi\right]$ and $\gamma\in\left[0,2\pi\right]$ represent the Eulerian angles in the local basis vector, as shown in Figure 8.
@@
In their original proposal, the artery model was constructed based on the fully incompressible assumption, and the potential energy adopts $\tilde{I}_{4}:=J^{-2/3}I_{4}$ and $\tilde{I}_{6}:=J^{-2/3}I_{6}$ instead of $I_{4}$ and $I_{6}$ in the definition [26]. However, that model was identified to be plagued by pathological material behavior when being extended to compressible regimes. A modification was introduced in [56] to rectify the issue and the rectified energy form is given by (2.5).
@@
The dispersion parameter $\kappa$ characterizes the distribution of the fibers. When $\kappa=0$, there is no dispersion, and the model (2.5) recovers its predecessor [34], which is sometimes known as the HGO model. When $\kappa=1/3$, the fibers distribute equally in all directions and the model becomes isotropic and takes the exponential form, rendering it similar to the one of Demiray [22].
@@ 2.4 Tissue prestressing and an initialization procedure
In patient-specific modeling, the vascular wall is imaged in a state under physiological loading. The imaged configuration is often taken as the referential or initial configuration, but it is not the natural configuration for the wall. This means that there is a stress field existing in the imaged configuration. The stress field in the imaged configuration is referred to as prestress and denoted by $\bm{\sigma}_{0}$. This prestress balances the in vivo blood pressure and WSS.
We incorporate the prestress term $\bm{\sigma}_{0}$ into the Cauchy stress to furnish a modified constitutive relation $\bm{\sigma}=\bm{\sigma}_{\mathrm{d}}-p\bm{I}+\bm{\sigma}_{0}$. To determine the value of $\bm{\sigma}_{0}$ at each quadrature point, we perform a fixed-point iteration so that the wall does not exhibit deformation under the loading of the fluid. A rigid-walled simulation is first conducted to acquire the physiological pressure and WSS distributions at diastole. We mention that the patient-specific image data investigated in this work is indeed cardiac-gated and synchronized to the cardiac cycle, which guarantees the compatibility of the geometric model of the wall and the prestress to be determined. Upon attaining a steady-state flow state, the solution is loaded as the input data to generate the prestress $\bm{\sigma}_{0}$. The algorithm is detailed below, drawing inspiration from a prior work grounded in pure displacement formulation [39].
@@ 3 Results
We begin by evaluating the local basis vectors generated using the method presented in Section 2.2 to demonstrate the effectiveness of the proposed morphology-based strategy. After that, two vascular FSI studies are considered, focusing on the impact of the fiber-reinforced model on the prediction of biomechanical factors. Our FSI model can be summarized as follows.
@@
    The FSI formulation is based on a unified continuum formulation, which allows a convenient way of monolithic coupling [48, 70]. The mesh motion in the fluid subdomain is governed by the harmonic extension algorithm.
@@
    Blood is modeled as a Newtonian fluid.
@@
    The variational multiscale formulation is employed as the numerical model to provide pressure stabilization for both fluids and solids and to facilitate large-eddy simulation of complex blood flows [48, 49]. In particular, linear tetrahedral element is used for the interpolation of all field variables.
@@
  5. 5.
    In the nonlinear solver, the relative tolerance is set to $1\times 10^{-6}$ as the stopping criterion, while the relative tolerance for solving the mesh motion equations is set to $5\times 10^{-4}$.
@@ 3.1 Evaluation of methods for generating local basis vectors
In this section, we performed a comparison between the approach for generating local basis vectors presented in Section 2.2 with the method described in [4, 52]. The latter employs static linear elastic analysis to generate local basis vectors. In specific, the geometry was loaded with internal pressure, and a Robin boundary condition is applied on the exterior surface to prevent bending and rigid motion. The principal directions of maximum, mid, and minimum principal stresses define the local circumferential, axial, and radial directions of the vessel, respectively.
The comparison was conducted using a T-shaped geometry. The local circumferential and axial directions based on the elastic stresses are illustrated in Figure 9 (a) and (b), respectively. In Figure 9 (c) and (d), we present the local circumferential and axial vectors obtained using the proposed method. Both methods yield satisfactory local circumferential directions with minor differences. Nevertheless, for the local radial direction near the bifurcation region, the direction of the mid-principal stress noticeably differs from the actual axial direction. In contrast, our method robustly delivers higher-quality definition of the local axial direction.
@@
Figure 12: Displacement of the arterial wall at $t=0.536$ s.
Figure 13: Von Mises stress distribution of the arterial wall at $t=0.536$ s.
@@ 3.2 Curved tube FSI benchmark
We consider the benchmark proposed in [8] to investigate the fiber reinforced material model in the setting of FSI. The geometric setting is illustrated in Figure 10, and it represents an idealized coronary artery. The axial displacement components of all nodes on the inlet and outlet annular surfaces are fixed, while the nodes marked by the green line are only allowed to move in the radial direction. On the inlet surface, the fluid velocity is prescribed with a parabolic profile with the flow rate over time shown in Figure 10. In this example, we do not have the prestress field initialized for the tissue. Instead, the curved elastic tube is inflated with a physiological pressure applied on the interior surface. This is referred to as the prestretch procedure [8]. To have the tube inflated in a robust manner, the flow rate increases from zero to $3.0$ mL/s in $0.1$ s using a cosine-type ramp function of time and maintained at that value for another $0.1$ s. The chosen cosine-type ramp function assists avoiding undesirable oscillations [8]. On the outlet surface of the fluid subdomain, a resistance boundary condition is applied with the resistance value being $35553.33$ $\cdots$/cm5 and zero distal pressure, which gives a pressure of $80$ mmHg at the flow rate of $3.0$ mL/s. At time $t=0.2$ s, the flow and elastic tube can be viewed to reach a steady state, and the pulsatile flow profile is turned on. As is shown in Figure 10, the maximum flow rate reaches $7.97$ mL/s at time $t=0.536$ s, and the cardiac cycle is $1.0$ s. This pulsatile flow profile mimics that of a right coronary artery. On the outlet of the fluid subdomain, a resistance boundary condition is applied with the resistance and distal pressure values being $7720.86$ $\cdots$/cm5 and $83472.39$ dyn/cm2, which gives the outlet pressure of $80$ mmHg at $t=0.2$ and $108$ mmHg at $t=0.536$. The fluid density and dynamic viscosity are set to be 1.0 g/cm3 and 0.04 dyn·s/cm2, respectively. In the numerical study, we use a mesh with 517800 linear tetrahedral elements, consisting of 336000 solid elements and 181800 fluid elements. Time integration are performed with a fixed time step size $\Delta t=1.0\times 10^{-4}$ s.
@@ 3.2.1 The curved tube fluid–structure benchmark problem
We first investigate the impact of the constitutive models of the elastic wall, and the results are compared with the results reported in [9] for verification purposes. Three different material models are considered:
@@
The material parameters for the GOH model are fitted based on a uniaxial tension test of the arterial media layer [14, 33], while the shear moduli for Cases 1 and 2 are chosen to approximate the stiffness of the experimental data at sufficiently large stretches [9, Figure 1]. For the neo-Hookean model with even lower shear modulus, we did observe non-robust numerical behavior, consistent with the observation reported in [9].
@@
Figure 11 illustrates the flow rate, area, and averaged pressure over time calculated on the outflow surface. The results matched well with those of [8]. The flow rate and averaged pressure of these three models are very close. This is expected because the former is dictated by the global mass conservation, and the latter follows the imposed boundary condition. The area changes or the displacement on the outlet annular surface are larger for the neo-Hookean models, likely due to the lack of fiber reinforcement.
@@
Figures 12 and 13 depicts the displacement and von Mises stress of the arterial wall at time $t=0.536$ s, respectively. The black grid in Figure 12 represents the initial position of the elastic tube. A significant difference between the isotropic and anisotropic models can be clearly observable, while the differences between the two isotropic models are difficult to discern. The displacement magnitude of the neo-Hookean models are significantly smaller. In the meantime, the neo-Hookean tube bended upward near the outlet, while in the GOH bended medially near the inlet. The anisotropic model exhibit a localized stress distribution, suggesting the fibers are sufficiently stretched near the the luminal interface and carry a significant amount of loads.
3.2.2 Parameter test of GOH material model
In this section, the GOH model is further investigated by inspecting the impact of the fiber orientation and dispersion. The problem settings remain the same as those in Section 3.2.1. The values of $\theta$ are taken as $40^{\circ}$, $42^{\circ}$, $44^{\circ}$, $46^{\circ}$, $48^{\circ}$, and $50^{\circ}$, while the values of $\kappa$ are set to $0.0$, $0.111$, $0.222$, and $0.333$. In particular, when $\kappa=0.0$, there is no fiber dispersion and the model degenerates into the predecessor of the GOH model [34], and when $\kappa=0.333$, the fiber are distributed equally in all directions, resulting in an isotropic model with an exponential form.
The changes of the area over time at the outlet are shown in Figure 14. It can be observed that under a fixed value of $\kappa$, the area change becomes more pronounced when the angle $\theta$ is smaller, meaning the fibers are more aligned with the circumferential direction. In the considered FSI problem, the deformation behavior of the vessel wall predominantly involves radial expansion caused by the flow pressure. Therefore, when the fiber orientation is closer to the circumferential direction, it more effectively activates fiber reinforcement. Moreover, the results show that the differences between the curves of different angle values diminish as $\kappa$ increases. This is because larger $\kappa$ values indicate a more dispersed fiber orientation, rendering the impact of $\theta$ less significant. In particular, when $\kappa=0.333$, the material becomes isotropic and the angle $\theta$ no longer affect the material behavior.
Figure 15 shows the displacement distribution of the wall at time $t=0.536$ s, with the black grid representing the initial position of the wall. The results of $\theta=40^{\circ}$ and $\theta=50^{\circ}$ are depicted. When comparing the results of different angle values with the same $\kappa$ value, the displacements of $\theta=40^{\circ}$ are consistently larger than those of $\theta=50^{\circ}$. This is because the fibers oriented with $\theta=50^{\circ}$ provides more reinforcement of the wall. When comparing the results of different $\kappa$ values with the same angle, the wall displacement decreases with the increase of the fiber dispersion. Interestingly, the deformation of the wall becomes more similar to that of the neo-Hookean model when $\kappa$ gets closer to $0.333$. It can be observed that wall bended upward near the outlet and the displacement becomes negligible near the inlet when $\kappa=0.333$. Conversely, when $\kappa=0.0$, the wall bended medially near the inlet with almost no displacement at the outlet. This is consistent with the prior observation made in [9] and Figure 12, suggesting that the anisotropic and isotropic material models behaves substantially different under physiological loading.
Figure 16 further illustrates the von Mises stress distribution at time $t=0.536$ s. When the angle $\theta=50^{\circ}$, the von Mises stress increases with the increase of $\kappa$ value. However, when $\theta=40^{\circ}$, the relation between the von Mises stress and the dispersion parameter is less straightforward. The stress initially decreases and then increases as $\kappa$ increases from $0.0$ to $0.333$. To gain an understanding of this phenomenon, we monitored the values of $\mathcal{I}:=\kappa{I}_{1}+(1-3\kappa){I}_{4}-1$, as shown in Figure 17. The term $\mathcal{I}$ appears in the exponential function of the strain energy (2.5), and its variation pattern with respect to $\kappa$ is mirrors that of the von Mises stress for both angles. The value of $\mathcal{I}$ is influenced by the combined effects of the dispersion parameter and the deformation state, and this partly explains the von Mises stresses illustrated in Figure 16.
@@ 3.2.2 Parameter test of GOH material model
3.3 Patient-specific study
In this example, we consider an image-based abdominal aorta (AA) model, which includes multiple visceral arteries (hepatic, splenic, renal, and superior mesenteric arteries (SMA)). The geometric model was constructed based on the MRI scan data from a healthy adult male [57, 65], employing the approach outlined in Section 2.1 (Figure 18 (a)). The respiratory-gated, non-contrast-enhanced magnetic resonance angiogram images were obtained using the inflow inversion recovery sequence on a 3.0 T magnetic resonance scanner (Discovery MR750, GE Medical Systems) equipped with a 32-channel coil. More details of this scanner can be found in [65].

### v2

@@ 2.3 Constitutive model for arterial wall
In this section, we introduce the arterial wall constitutive models. The arterial wall is modeled as a slightly compressible or fully incompressible hyperelastic material. Given the referential configuration $\Omega_{0}\subset\mathbb{R}^{3}$, its motion at time $t$ is given by an orientation-preserving one-to-one mapping $\bm{\varphi}_{t}(\cdot)$ that maps the body to the current configuration $\Omega_{t}=\bm{\varphi}_{t}(\Omega_{0})$. The kinematics is locally described by the deformation gradient $\bm{F}:=\partial\bm{\varphi}_{t}/\partial\bm{X}$ for $\bm{X}\in\Omega_{0}$. Its determinant is known as Jacobian and denoted by $J$. The right Cauchy-Green deformation tensor is given by $\bm{C}:=\bm{F}^{T}\bm{F}$. On the referential configuration, a scalar field $P$ is introduced to represent the thermodynamic pressure. The material behavior is characterized by the free energy of the form,
@@
where $G$ represents the Gibbs free energy. It is a scalar-valued function and is additively decomposed into the contributions of $G_{\mathrm{d}}$ and $G_{\mathrm{v}}$. The deformation is often decoupled into isochoric and volumetric parts following the decomposition of the Flory type, i.e., $\bm{F}=J^{1/3}\tilde{\bm{F}}$, and correspondingly the unimodular right Cauchy-Green deformation tensor is defined as $\tilde{\bm{C}}:=\tilde{\bm{F}}^{T}\tilde{\bm{F}}$. If the energy function $G_{\mathrm{a}}$ depends on $\tilde{\bm{C}}$, it can be interpreted as the energy associated with isochoric deformation, and $G_{\mathrm{b}}$ is then the bulk or volumetric free energy. Through a Coleman-Noll-type analysis [48], we may obtain the constitutive relations for the second Piola-Kirchhoff stress $\bm{S}$, the Cauchy stress $\bm{\sigma}$, and the density $\rho$ based on the free energy $G$ as
@@
In the theory constructed based on the Gibbs free energy, the pressure $P$ acts as a primitive variable, rendering a mixed formulation suitable for (nearly-)incompressible materials [48]. If the energy is taken simply as $G_{\mathrm{b}}=P$, the one has $\rho=\rho_{0}$ and the material is fully incompressible. In this work, we adopt the following form for the volumetric energy,
@@
where $K$ is the bulk modulus. This volumetric energy is related to the volumetric energy $K(J^{2}-1-2\ln(J))/4$ through a Legendre transformation [48]. A specific functional form of $G_{\mathrm{d}}$ completes the definition of the constitutive model.
@@
where $\tilde{I}_{1}:=\mathrm{tr}\tilde{\bm{C}}$ is the first principal invariant of $\tilde{\bm{C}}$, and the parameter $\mu$ is the shear modulus. This model enjoys a compact energy form with a single material parameter and can be rigorously derived based on the microscopic structure using the Gaussian distribution. Despite its wide usage in vascular FSI analysis [11, 49, 69], we need to point out that it was originally designed for vulcanized rubber and its predictive capability is limited to moderate deformation states [80].
@@
and the latter two describe the stretch along the fiber directions. The parameters $\mu$ and $k_{1}$ are stress‐like, representing the stiffness of the ground matrix and collagen fibers; $k_{2}$ is a dimensionless parameter; the unit vectors $\bm{a}_{0}$ and $\bm{b}_{0}$ represent the mean orientation of the embedded fibers; the parameter $\kappa\in[0,1/3]$ characterizes the fiber dispersion. The first term in (2.5) represents the potential energy due to the ground matrix, which is treated as an isotropic neo-Hookean material; the exponential terms in (2.5) represent the anisotropic free energy contributions due to the two families of fibers. In this work, we consider the two families of fibers disposed in a symmetric manner with respect to the axial direction. Therefore, the fiber directions $\mathbf{a}_{0}(\bm{X})$ and $\mathbf{b}_{0}(\bm{X})$ at a material point $\bm{X}\in\Omega_{0}$ can be defined by the local basis vector $\{\bm{e}_{\mathrm{r}}(\bm{X}),\bm{e}_{\mathrm{c}}(\bm{X}),\bm{e}_{\mathrm{a}}(\bm{X})\}$ as follow,
@@
where $\theta\in\left[0,\pi\right]$ and $\gamma\in\left[0,2\pi\right]$ are the Eulerian angles, as shown in Figure 8.
@@
In their original proposal, the artery model was constructed based on the fully incompressible assumption, and the potential energy adopts $\tilde{I}_{4}:=J^{-2/3}I_{4}$ and $\tilde{I}_{6}:=J^{-2/3}I_{6}$ instead of $I_{4}$ and $I_{6}$ in the definition [26]. However, that model was plagued by pathological material behavior when being extended to compressible regimes. A modification was introduced in [56] to rectify the issue and the rectified energy is given by (2.5).
@@
The dispersion parameter $\kappa$ characterizes the distribution of the fibers. When $\kappa=0$, there is no dispersion, and the model (2.5) recovers its predecessor [34], which is known as the HGO model. When $\kappa=1/3$, the fibers distribute equally in all directions and the model becomes isotropic and takes the exponential form, rendering it similar to the one of Demiray [22].
@@ 2.4 Tissue prestressing and an initialization procedure
In patient-specific modeling, the vascular wall is imaged in a state under physiological loading, meaning that there is a stress field in the imaged configuration. The stress field is often referred to as prestress and denoted by $\bm{\sigma}_{0}$. This prestress balances the in vivo blood pressure and WSS. We incorporate the prestress term $\bm{\sigma}_{0}$ into the Cauchy stress to furnish a modified constitutive relation $\bm{\sigma}=\bm{\sigma}_{\mathrm{d}}-p\bm{I}+\bm{\sigma}_{0}$. To determine the value of $\bm{\sigma}_{0}$ at each quadrature point, we perform a fixed-point iteration so that the wall does not exhibit deformation under the loading of the fluid. A rigid-walled simulation is first conducted to acquire the physiological pressure and WSS distributions. We mention that the patient-specific image data investigated in this work is indeed cardiac-gated and synchronized to the cardiac cycle, which guarantees the compatibility of the geometric model of the wall and the prestress to be determined. Upon attaining a steady-state flow state, the solution is loaded as the input data to generate the prestress $\bm{\sigma}_{0}$. The algorithm is detailed below, drawing inspiration from a prior work grounded in pure displacement formulation [39].
@@ 3 Results
We begin by evaluating the local basis vectors generated using the method presented in Section 2.2 to demonstrate the effectiveness of the proposed morphology-based strategy. After that, two vascular FSI studies are considered, focusing on the impact of the fiber-reinforced model on the prediction of biomechanical factors. The key features of our FSI model can be summarized as follows.
@@
    The formulation is based on the unified continuum formulation, allowing a convenient way of monolithic coupling [48, 70]. The mesh motion in the fluid subdomain is governed by the harmonic extension algorithm.
@@
    The blood is modeled as a Newtonian fluid.
@@
    The variational multiscale formulation is employed as the numerical model to provide pressure stabilization for both fluids and solids and to facilitate large-eddy simulation of complex blood flows [48, 49].
@@
    Linear tetrahedral element is used for the interpolation of all field variables.
  5. 5.
@@ 3.1 Evaluation of methods for generating local basis vectors
In this section, we perform a comparison between the approach for generating local basis vectors presented in Section 2.2 with the method described in [4, 52]. The latter employs static linear elastic analysis to generate local basis vectors. In specific, the geometry is loaded with internal pressure, and a Robin boundary condition is applied on the exterior surface to prevent bending and rigid motion. The principal directions of the maximum, mid, and minimum principal stresses define the local circumferential, axial, and radial directions of the vessel, respectively.
The comparison is conducted using a T-shaped geometry. The local circumferential and axial directions based on the elastic stresses are illustrated in Figure 9 (a) and (c), respectively. In Figure 9 (b) and (d), we present the local circumferential and axial vectors obtained using the proposed method. Both methods yield satisfactory local circumferential directions with minor differences. Nevertheless, for the local radial direction near the bifurcation region, the direction of the mid-principal stress noticeably differs from the actual axial direction. In contrast, our morphology-based approach robustly delivers higher-quality definition of the local axial direction.
@@
Figure 12: The deformation of the arterial wall at $t=0.536$ s.
Figure 13: The distribution of the von Mises stress in the arterial wall at $t=0.536$ s.
@@ 3.2 Curved tube FSI benchmark
We consider the benchmark proposed in [8, 9] to investigate the fiber reinforced material model in the setting of FSI. The geometric setting is illustrated in Figure 10, representing an idealized coronary artery. The axial displacement components of all nodes on the inlet and outlet annular surfaces are fixed, and the nodes marked by the green line are only allowed to move in the radial direction. On the inlet surface, the fluid velocity is prescribed with a parabolic profile with the flow rate over time shown in Figure 10. In this example, we do not have the prestress field initialized for the tissue. Instead, the curved elastic tube is inflated with a physiological pressure applied on the interior surface. This is referred to as the prestretch procedure [8]. To have the tube inflated in a robust manner, the flow rate increases from zero to $3.0$ mL/s in $0.1$ s using a cosine-type ramp function of time and maintains at that value for another $0.1$ s. The chosen cosine-type ramp function assists in avoiding undesirable oscillations [8]. On the outlet surface of the fluid subdomain, a resistance boundary condition is applied with the resistance value being $35553.33$ $\cdots$/cm5 and zero distal pressure, which gives a pressure of $80$ mmHg at the flow rate of $3.0$ mL/s. At time $t=0.2$ s, the flow and elastic tube can be viewed to reach a steady state, and the pulsatile flow profile is turned on. As is shown in Figure 10, the maximum flow rate reaches $7.97$ mL/s at time $t=0.536$ s, and the cardiac cycle is $1.0$ s. This pulsatile flow profile mimics that of a right coronary artery. At the outlet of the fluid subdomain, a resistance boundary condition is applied with the resistance and distal pressure values being $7720.86$ $\cdots$/cm5 and $83472.39$ dyn/cm2, which gives the outlet pressure of $80$ mmHg at $t=0.2$ and $108$ mmHg at $t=0.536$. The fluid density and dynamic viscosity are set to be 1.0 g/cm3 and 0.04 dyn·s/cm2, respectively. In the numerical study, we use a mesh with $5.18\times 10^{5}$ linear tetrahedral elements, consisting of $3.36\times 10^{5}$ solid elements and $1.82\times 10^{5}$ fluid elements. Time integration is performed with a fixed time step size $\Delta t=1.0\times 10^{-4}$ s. We mention that a mesh independence study has been performed, and the reported mesh and time step sizes are the finest spatiotemporal resolution we used.
@@ 3.2.1 The curved tube fluid–structure benchmark problem
We first investigate the impact of the constitutive models of the elastic wall, and the results are compared with those reported in [9] for verification purposes. Three different material models are considered:
@@
The material parameters for the GOH model are fitted based on a uniaxial tension test of the arterial media layer [14, 33], while the shear moduli for Cases 1 and 2 are chosen to approximate the stiffness of the experimental data at sufficiently large stretches [9, Figure 1]. For the neo-Hookean model with even lower shear modulus, we do observe non-robust numerical behavior, consistent with the observation reported in [9].
@@
Figure 11 illustrates the flow rate, area, and averaged pressure over time calculated on the outflow surface. The results match well with those of [8]. The flow rate and averaged pressure of these three models are very close. This is expected because the former is dictated by the global mass conservation, and the latter follows the imposed boundary condition. The area changes or the displacement on the outlet annular surface are larger for the neo-Hookean models, possibly due to the lack of fiber reinforcement.
@@
Figures 12 and 13 depict the displacement and von Mises stress of the arterial wall at time $t=0.536$ s, respectively. The black grid in Figure 12 represents the initial position of the elastic tube. A significant difference between the isotropic and anisotropic models is clearly observable, while the differences between the two isotropic models are difficult to discern. The displacement magnitude of the neo-Hookean models is significantly smaller. In the meantime, the neo-Hookean tube bends upward near the outlet, while the GOH bends medially near the inlet. The anisotropic model exhibits a localized stress distribution, suggesting the fibers are sufficiently stretched near the luminal interface and carry a significant amount of loads.
@@
3.2.2 Parameter test of GOH material model
In this section, the GOH model is further investigated by inspecting the impact of the fiber orientation and dispersion. The problem settings remain the same as those in Section 3.2.1. The values of $\theta$ are taken as $40^{\circ}$, $42^{\circ}$, $44^{\circ}$, $46^{\circ}$, $48^{\circ}$, and $50^{\circ}$, while the values of $\kappa$ are set to $0.0$, $0.111$, $0.222$, and $0.333$.
The changes in the area over time at the outlet are shown in Figure 14. It can be observed that under a fixed value of $\kappa$, the area change becomes more pronounced when the angle $\theta$ is smaller, meaning the fibers are more aligned with the axial direction. In the considered FSI problem, the deformation behavior of the vessel wall predominantly involves radial expansion caused by the flow pressure. Therefore, when the fiber orientation is closer to the circumferential direction, it more effectively activates fiber reinforcement. Moreover, the results show that the differences between the curves of different angle values diminish as $\kappa$ increases. This is because larger $\kappa$ values indicate a more dispersed fiber orientation, rendering the impact of $\theta$ less significant. In particular, when $\kappa=0.333$, the material becomes isotropic and the angle $\theta$ no longer affects the material behavior.
Figure 15 shows the displacement distribution of the wall at time $t=0.536$ s, with the black grid representing the initial position of the wall. The results of $\theta=40^{\circ}$ and $\theta=50^{\circ}$ are depicted. When comparing the results of different angle values with the same $\kappa$ value, the displacements of $\theta=40^{\circ}$ are consistently larger than those of $\theta=50^{\circ}$. This is because the fibers oriented with $\theta=50^{\circ}$ provide more reinforcement of the wall. When comparing the results of different $\kappa$ values with the same angle, the wall displacement decreases with the increase of the fiber dispersion. Interestingly, the deformation of the wall becomes more similar to that of the neo-Hookean model when $\kappa$ gets closer to $0.333$. It can be observed that wall bends upward near the outlet and the displacement becomes negligible near the inlet when $\kappa=0.333$. Conversely, when $\kappa=0.0$, the wall bends medially near the inlet with almost no displacement at the outlet. This is consistent with the prior observation made in [9] and Figure 12, indicating that the deformation state of Case 3 in Figure 12 is in fact caused by the anisotropy rather than the exponential energy form. The differences in the deformation states suggest that considering anisotropy in the tissue model may lead to substantially different behavior under physiological loading.
Figure 16 further illustrates the von Mises stress distribution at time $t=0.536$ s. When the angle $\theta=50^{\circ}$, the von Mises stress increases with the increase of $\kappa$ value. However, when $\theta=40^{\circ}$, the relation between the von Mises stress and the dispersion parameter is less straightforward. The stress initially decreases and then increases as $\kappa$ increases from $0.0$ to $0.333$. To gain an understanding of this phenomenon, we monitor the values of $\mathcal{I}:=\kappa{I}_{1}+(1-3\kappa){I}_{4}-1$, as shown in Figure 17. The term $\mathcal{I}$ appears in the exponential function of the strain energy (2.5), and its variation pattern with respect to $\kappa$ mirrors that of the von Mises stress for both angles. The value of $\mathcal{I}$ is influenced by the combined effects of the dispersion parameter and the deformation state, and this explains the von Mises stresses illustrated in Figure 16.
3.3 Patient-specific abdominal aorta model
In this example, we consider an image-based abdominal aorta (AA) model, which includes multiple visceral arteries (hepatic, splenic, renal, and superior mesenteric arteries (SMA)). The geometric model is constructed based on the MRI data from a healthy adult male [57, 65] (Figure 18 (a)). The non-contrast-enhanced magnetic resonance angiogram images are obtained using the inflow inversion recovery sequence on a 3.0 T magnetic resonance scanner (Discovery MR750, GE Medical Systems) equipped with a 32-channel coil. More details of this scanner can be found in [65].